OpenStudy (anonymous):

what is the equation of the vertical asymptotes of the graph g(x)=-log(x)+2

OpenStudy (amistre64):

when is log(x) undefined?

OpenStudy (anonymous):

what?

OpenStudy (amistre64):

.... for what values of x, is the log function .... not defined?

OpenStudy (anonymous):

f(x)=log(x)

OpenStudy (amistre64):

do you recall what a vertical asymptote is?

OpenStudy (anonymous):

i dont think so

OpenStudy (amistre64):

it should be presented to you in your material ...

OpenStudy (amistre64):

if you lok at a graph of the function; it is related to all the values of x that cannot be defined for the function: like 1/x has a vertical asymptote at x=0, since 1/0 cannot be defined

OpenStudy (anonymous):

ill just write it all out for you. describe the transformations on the following graph of f(x)=log(x). state the placement of the vertical symptote and x intercept after the transformation. for example, vertical pellet up 2 or reflected about the x-axis are descriptions. g(x)=-log(x)+2

OpenStudy (amistre64):

it looks like they expect you to have familiarity with the log fucntion in order to address the problem to start with

OpenStudy (amistre64):

log(x) has an x intercept at x=1, and a vertical asymptote at x=0 how does the shift alter these properties?

OpenStudy (anonymous):

i already have the transformation, its the graph of G(x)=-log(x)+2 reflects across the x axis and upward 5.

OpenStudy (amistre64):

+2 is not upward by 5 ...

OpenStudy (anonymous):

yah my mistake

OpenStudy (amistre64):

g = -logx + 2 0 = b^(-logx + 2) b^0 = b^(-logx) * b^2 1/b^2 = b^(log(1/x) 1/b^2 = 1/x x = b^2

OpenStudy (amistre64):

in some courses, log is generaic and represents the natural log; in other courses log is a base10 construction

OpenStudy (anonymous):

b represents what?

OpenStudy (amistre64):

b = base that is consistent with whatever the course defines log to be :)

OpenStudy (anonymous):

so the answer is b^2

OpenStudy (amistre64):

whatever the base is; the solution would be the square of the base .. yes

OpenStudy (amistre64):

0 = 2 - log(x), when x = b^2 0 = 2 - log(b^2) 0 = 2 - 2log(b) 0 = 2 - 2(1)

OpenStudy (anonymous):

for my previous function, g(x)=log(x-5) the vertical asymptote is 5

OpenStudy (amistre64):

that looks appropriate yes.

OpenStudy (anonymous):

so why is this function so complicated?

OpenStudy (amistre64):

the only complication is in the x intercept. they have not shifted the graph left or right; only up and down you therefore need to define the value of x when g(x)=0

OpenStudy (anonymous):

so whats the x intrercept of the function

OpenStudy (amistre64):

in general; x = (vertical asymptote) is affected by a left to right shift of +c like this: x+c = vertical asymptote x = vertical asymptote - c

OpenStudy (amistre64):

i already covered how to determine the x intercept of this function ...

OpenStudy (anonymous):

ok i think the x intercept is (0,2)

OpenStudy (amistre64):

it wil be (0,b^2) ; so unless your log base is in sqrt2 .... you might want to reconsider that